# Strichartz estimates for the generalized Zakharov-Kuznetsov equation on ×$\mathbb{R}\times\mathbb{T}$ and applications

### v1

@@ 4.2. Improved LWP for mZK
$\lvert 3(\xi_{0}+\xi_{2})^{2}-(q_{0}+q_{2})^{2}\rvert\ll\lvert\xi_{0}+\xi_{2}\rvert\sim\lvert\xi_{0}\rvert.$
Subtracting the two estimates yields $\lvert 3\xi_{0}\xi_{2}-q_{0}q_{2}\rvert\ll\lvert\xi_{0}\rvert$, hence
$\Big{\lvert}3\lvert\xi_{2}\rvert-\frac{\lvert q_{0}\rvert\lvert q_{2}\rvert}{\lvert\xi_{0}\rvert}\rvert\ll 1$, and from the two assumptions (ii.3) and (ii.3.1), it follows that $\lvert\xi_{2}\rvert\ll\lvert q_{2}\rvert$. Altogether, we thus obtain $\lvert q_{0}\rvert\ll\lvert\xi_{0}\rvert$, and taking into account once again the general assumption (ii.3) and the fact that $\lvert\xi_{0}-\xi_{3}\rvert\ll\lvert\xi_{0}\rvert$, it follows that $\lvert q_{3}\rvert\ll\lvert\xi_{3}\rvert$ as well. From $\lvert q_{0}\rvert\ll\lvert\xi_{0}\rvert$ and $\lvert q_{3}\rvert\ll\lvert\xi_{3}\rvert$, it now follows that $\lvert 3(\xi_{0}+\xi_{3})^{2}-(q_{0}+q_{3})^{2}\rvert\gtrsim\lvert\xi_{0}+\xi_{3}\rvert^{2}\geq\lvert\xi_{0}+\xi_{3}\rvert$, which means that we can resort to variant (33) of the bilinear refinement (30) and proceed exactly as in the case $\lvert 3(\xi_{0}+\xi_{2})^{2}-(q_{0}+q_{2})^{2}\rvert\gtrsim\lvert\xi_{0}+\xi_{2}\rvert$. This concludes the discussion of this subcase.
(ii.3.2.1) $\lvert\xi_{\text{med}}+\xi_{\text{max}}\rvert\gtrsim\lvert\xi_{\text{med}}\rvert$:
Under this assumption, the estimates
$\lvert\xi_{0}\rvert\lesssim\lvert\xi_{0}+\xi_{1}\rvert$, $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{2}+\xi_{3}\rvert$, $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{0}+\xi_{3}\rvert$, and $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{1}+\xi_{2}\rvert$ hold. If $\lvert 3(\xi_{i}+\xi_{j})^{2}-(q_{i}+q_{j})^{2}\rvert\gtrsim\lvert\xi_{i}+\xi_{j}\rvert$ holds for a tuple $(i,j)\in\{(0,1),(2,3),(0,3),(1,2)\}$, then we may argue exactly as in case (ii.3.1), so that variants of (2) and (30) lead to the desired multilinear estimate (42), valid for every $s>\frac{3}{8}$. Consequently, we may assume
$\lvert 3(\xi_{0}+\xi_{1})^{2}-(q_{0}+q_{1})^{2}\rvert\ll\lvert\xi_{0}+\xi_{1}\rvert\sim\lvert\xi_{0}\rvert,$
@@
$\lvert 3(\xi_{0}+\xi_{3})^{2}-(q_{0}+q_{3})^{2}\rvert\ll\lvert\xi_{0}+\xi_{3}\rvert\sim\lvert\xi_{0}\rvert,$
and
$\lvert 3(\xi_{1}+\xi_{2})^{2}-(q_{1}+q_{2})^{2}\rvert\ll\lvert\xi_{1}+\xi_{2}\rvert\sim\lvert\xi_{0}\rvert.$
Subtracting the first estimate from the second, and the third from the fourth yields $\Big{\lvert}3\lvert\xi_{1}\rvert-\frac{\lvert q_{0}\rvert\lvert q_{1}\rvert}{\lvert\xi_{0}\rvert}\rvert\ll 1$ and $\Big{\lvert}3\lvert\xi_{3}\rvert-\frac{\lvert q_{0}\rvert\lvert q_{3}\rvert}{\lvert\xi_{0}\rvert}\rvert\ll 1$, respectively. From the combination of $\lvert\xi_{1}\rvert\ll\lvert q_{1}\rvert$ (note the assumptions (ii.3) and (ii.3.2)) and $\Big{\lvert}3\lvert\xi_{1}\rvert-\frac{\lvert q_{0}\rvert\lvert q_{1}\rvert}{\lvert\xi_{0}\rvert}\rvert\ll 1$, it then follows that $\lvert q_{0}\rvert\ll\lvert\xi_{0}\rvert$, and the latter, together with $\Big{\lvert}3\lvert\xi_{3}\rvert-\frac{\lvert q_{0}\rvert\lvert q_{3}\rvert}{\lvert\xi_{0}\rvert}\rvert\ll 1$, ultimately leads to $\lvert\xi_{3}\rvert\ll\lvert q_{3}\rvert$, which contradicts $\lvert\xi_{3}\rvert\sim\lvert\xi_{0}\rvert\sim\lvert(\xi_{0},q_{0})\rvert\sim\lvert(\xi_{3},q_{3})\rvert\geq\lvert q_{3}\rvert$. Thus, this subcase is also resolved.
(ii.3.2.2) $\lvert\xi_{\text{med}}+\xi_{\text{max}}\rvert\ll\lvert\xi_{\text{med}}\rvert$:
We have $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{0}+\xi_{2}\rvert$ and $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{1}+\xi_{3}\rvert$, so that the case $\lvert 3(\xi_{i}+\xi_{j})^{2}-(q_{i}+q_{j})^{2}\rvert\gtrsim\lvert\xi_{i}+\xi_{j}\rvert$ for some $(i,j)\in\{(0,2),(1,3)\}$, can be handled exactly as in (ii.3.1). If in both cases we cannot rely on the bilinear refinement (30), then we have
$\lvert 3(\xi_{0}+\xi_{2})^{2}-(q_{0}+q_{2})^{2}\rvert\ll\lvert\xi_{0}+\xi_{2}\rvert\sim\lvert\xi_{2}\rvert$
and
$\lvert 3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}\rvert\ll\lvert\xi_{1}+\xi_{3}\rvert\sim\lvert\xi_{2}\rvert,$
and substracting the two estimates yields $\Big{\lvert}3\lvert\xi_{0}\rvert-\frac{\lvert q_{0}\rvert\lvert q_{2}\rvert}{\lvert\xi_{2}\rvert}\rvert\ll 1$. Moreover, since $\lvert\xi_{0}\rvert\ll\lvert q_{0}\rvert$ and $\lvert\xi_{1}\rvert\ll\lvert q_{1}\rvert$ also hold by (ii.3), (ii.3.2) and (ii.3.2.2), it follows further that $\lvert q_{2}\rvert\ll\lvert\xi_{2}\rvert$. Combining this with the general assumptions (ii.3) and (ii.3.2.2), we also obtain $\lvert q_{3}\rvert\ll\lvert\xi_{3}\rvert$, and thus we arrive at the combined conclusion
$\lvert\xi_{0}\rvert\ll\lvert q_{0}\rvert,\ \lvert\xi_{1}\rvert\ll\lvert q_{1}\rvert,\ \lvert q_{2}\rvert\ll\lvert\xi_{2}\rvert,\ \lvert q_{3}\rvert\ll\lvert\xi_{3}\rvert.$ (43)
For the further analysis, we now consider the resonance function $R$. A straightforward computation shows that
$R=-6(\xi_{1}+\xi_{2})(\xi_{1}+\xi_{3})(\xi_{2}+\xi_{3})+\sum_{i=1}^{3}\xi_{i}(\lvert(\xi_{0},q_{0})\rvert^{2}-\lvert(\xi_{i},q_{i})\rvert^{2}).$ (44)
@@
If $\lvert\xi_{2}+\xi_{3}\rvert\gtrsim\lvert\xi_{0}\rvert$, then by applying the reverse triangle inequality and taking into account (ii.3), $\lvert\xi_{1}+\xi_{2}\rvert\sim\lvert\xi_{\text{max}}\rvert$, and $\lvert\xi_{1}+\xi_{3}\rvert\sim\lvert\xi_{\text{max}}\rvert$, it follows that
$\lvert R\rvert\gtrsim\lvert\xi_{0}\rvert\lvert\xi_{\text{max}}\rvert^{2}.$
@@
$\lvert\xi_{0}\rvert\lvert\xi_{\text{max}}\rvert^{2}\lesssim\max_{i=0}^{3}\langle\sigma_{i}\rangle,$
@@
$I_{f}$ $\lesssim\lVert J^{-\frac{2}{9}-}fJ^{\frac{1}{9}+}\tilde{u}_{2}J^{\frac{1}{9}+}\tilde{u}_{3}\rVert_{L_{txy}^{2}}\lVert u_{1}\rVert_{X_{s,\frac{1}{2}+\varepsilon}}$
$\leq\lVert J^{-\frac{2}{9}-}f\rVert_{L_{Txy}^{6-}}\lVert J^{\frac{1}{9}+}\tilde{u}_{2}\rVert_{L_{Txy}^{6+}}\lVert J^{\frac{1}{9}+}\tilde{u}_{3}\rVert_{L_{Txy}^{6+}}\lVert u_{1}\rVert_{X_{s,\frac{1}{2}+\varepsilon}}.$
@@
which is precisely estimate (42), valid for every $s>\frac{1}{3}$, provided $\varepsilon>0$ is chosen small enough. We have thus seen that in the case $\lvert\xi_{2}+\xi_{3}\rvert\gtrsim\lvert\xi_{0}\rvert$, the resonance function leads us to the desired result. Conversely, if $\lvert\xi_{2}+\xi_{3}\rvert\ll\lvert\xi_{0}\rvert$, then we have $\lvert\xi_{0}\rvert\lesssim\lvert\xi_{0}+\xi_{1}\rvert$, while the relations established in (43) lead us to
$\lvert 3(\xi_{0}+\xi_{1})^{2}-(q_{0}+q_{1})^{2}\rvert\gtrsim\lvert\xi_{0}+\xi_{1}\rvert$. The bilinear refinement is thus applicable and, just as in case (ii.3.1), yields $s>\frac{3}{8}$.

### v3

@@ 4.2. Improved LWP for mZK
and we will first consider why, under these assumptions, ${\lvert}\xi_{3}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}$ must hold: Under the assumption ${\lvert}\xi_{3}{\rvert}\ll{\lvert}(\xi_{1},q_{1}){\rvert}$, it follows - taking into account the active assumptions (ii.3) and (ii.3.1) - that
${\lvert}q_{1}{\rvert}\sim{\lvert}q_{2}{\rvert}\sim{\lvert}q_{3}{\rvert}\sim{\lvert}q_{0}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}$,
and the two constraints mentioned above, which arise from the bilinear refinement of the linear $L^{4}$-estimate, further impose the sign configurations $(q_{1},q_{2},q_{3},q_{0})=(\pm,\pm,\mp,\pm)$. However, we also have ${\lvert}\xi_{1}+\xi_{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}$ and ${\lvert}\xi_{0}-\xi_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}$, which, in conjunction with (ii.3) and the signs of the $q_{i}$, leads us to ${\lvert}q_{1}-q_{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}$, ${\lvert}q_{0}+q_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}$, and
$2{\lvert}q_{1}+q_{3}{\rvert}={\lvert}q_{1}-q_{2}+q_{0}+q_{3}{\rvert}\leq{\lvert}q_{1}-q_{2}{\rvert}+{\lvert}q_{0}+q_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}$,
in contradiction to ${\lvert}3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}{\rvert}\ll{\lvert}\xi_{1}+\xi_{3}{\rvert}\sim{\lvert}\xi_{0}{\rvert}$. Consequently, it follows that ${\lvert}\xi_{3}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}$ must be true. In this situation, two cases must be distinguished: If ${\lvert}q_{3}{\rvert}\gtrsim{\lvert}(\xi_{1},q_{1}){\rvert}$, then from
${\lvert}q_{1}^{2}-{\lvert}(\xi_{3},q_{3}){\rvert}^{2}{\rvert}\leq{\lvert}{\lvert}(\xi_{1},q_{1}){\rvert}^{2}-{\lvert}(\xi_{3},q_{3}){\rvert}^{2}{\rvert}+3\xi_{1}^{2}\ll\xi_{0}^{2},$
${\lvert}q_{2}^{2}-{\lvert}(\xi_{3},q_{3}){\rvert}^{2}{\rvert}\leq{\lvert}{\lvert}(\xi_{2},q_{2}){\rvert}^{2}-{\lvert}(\xi_{3},q_{3}){\rvert}^{2}{\rvert}+3\xi_{2}^{2}\ll\xi_{0}^{2},$
and the two constraints coming from the bilinear refinement, we again obtain the following two admissible sign configurations: $(q_{1},q_{2},q_{3},q_{0})=(\pm,\pm,\mp,\pm)$. At this stage, however, we can precisely as previously discussed, deduce that ${\lvert}q_{1}+q_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}$, which is in contradiction with ${\lvert}3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}{\rvert}\ll{\lvert}\xi_{1}+\xi_{3}{\rvert}\sim{\lvert}\xi_{0}{\rvert}$. Thus, the sole remaining case to be examined is ${\lvert}q_{3}{\rvert}\ll{\lvert}(\xi_{1},q_{1}){\rvert}$. In this case, the active assumptions (ii.3) and (ii.3.1) also yield ${\lvert}q_{0}{\lvert}\ll{\lvert}(\xi_{1},q_{1}){\rvert}$ and ${\lvert}q_{1}{\rvert}\sim{\lvert}q_{2}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}$, which altogether imply that $q_{1}$ and $q_{2}$ must have opposite signs. In order to take advantage of this fact, we consider the resonance function $R$. A straightforward computation shows that
$R=-6(\xi_{1}+\xi_{2})(\xi_{1}+\xi_{3})(\xi_{2}+\xi_{3})+\sum_{i=1}^{3}\xi_{i}({\lvert}(\xi_{0},q_{0}){\rvert}^{2}-{\lvert}(\xi_{i},q_{i}){\rvert}^{2}),$ (43)
@@
and if we now assume ${\lvert}\xi_{1}+\xi_{2}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{1}{6}}\sim{\lvert}\xi_{\text{max}}{\rvert}^{\frac{1}{6}}$, then by applying the reverse triangle inequality and taking into account (ii.3), ${\lvert}\xi_{1}+\xi_{3}{\rvert}\sim{\lvert}\xi_{\text{max}}{\rvert}$, and ${\lvert}\xi_{2}+\xi_{3}{\rvert}\sim{\lvert}\xi_{\text{max}}{\rvert}$, it follows that
${\lvert}R{\rvert}\gtrsim{\lvert}\xi_{\text{max}}{\rvert}^{\frac{13}{6}}.$
@@
${\lvert}\xi_{\text{max}}{\rvert}^{\frac{13}{6}}\lesssim\max_{i=0}^{3}\langle\sigma_{i}\rangle,$
@@
$I_{f}$ $\lesssim{\lVert}J^{-\frac{2}{9}-}fJ^{\frac{1}{9}+}u_{2}J^{\frac{1}{9}+}u_{3}{\rVert}_{L_{txy}^{2}}{\lVert}u_{1}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}}$
$\leq{\lVert}J^{-\frac{2}{9}-}f{\rVert}_{L_{Txy}^{6-}}{\lVert}J^{\frac{1}{9}+}u_{2}{\rVert}_{L_{Txy}^{6+}}{\lVert}J^{\frac{1}{9}+}u_{3}{\rVert}_{L_{Txy}^{6+}}{\lVert}u_{1}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}}.$
@@
which is precisely the estimate (42), valid for every $s>\frac{1}{3}$, provided $\varepsilon>0$ is chosen small enough. We have thus seen that in the case ${\lvert}\xi_{1}+\xi_{2}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{1}{6}}$, the resonance function leads us to the desired result. Conversely, assume that ${\lvert}\xi_{1}+\xi_{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{6}}$. From (ii.3) and the fact that $q_{1}$ and $q_{2}$ have opposite signs, it then follows that ${\lvert}q_{1}+q_{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{6}}$, and we now aim to exploit this smallness condition. For dyadic $N\in 2^{\mathbb{N}_{0}}$ and $\beta\in\mathbb{R}$, we define the Fourier projectors $Q_{N}^{(\beta)}$ by setting
If we now set
$u_{1}=P_{N_{1}}u_{1},\ u_{2}=P_{N_{2}}u_{2},\ \text{and}\ u_{3}=P_{N_{3}}u_{3}$
for dyadic numbers $N_{1},N_{2},N_{3}\in 2^{\mathbb{N}_{0}}$, it follows that ${\lvert}q_{1}+q_{2}{\rvert}\ll N_{1}^{\frac{1}{6}}$, and redistributing derivatives gives [^6] We assume without loss of generality that ${\lvert}\xi_{1}{\rvert},{\lvert}\xi_{2}{\rvert}\gtrsim 1$. Otherwise, one may invoke the Sobolev embedding in $x$ without any loss of derivatives, which allows for the conclusion ${\lVert}u_{i}{\rVert}_{L_{Tx}^{4}L_{y}^{2}}\lesssim_{T}{\lVert}u_{i}{\rVert}_{X_{0,\frac{1}{2}+}}$, $i\in\{1,2\}$.
${\lvert}\xi_{0}{\rvert}\langle(\xi_{0},q_{0})\rangle^{s}\lesssim N_{1}^{\frac{3}{4}-2s}{\lvert}\xi_{0}{\rvert}^{\frac{1}{8}}{\lvert}\xi_{1}{\rvert}^{\frac{1}{8}}\langle(\xi_{1},q_{1})\rangle^{s}{\lvert}\xi_{2}{\rvert}^{\frac{1}{8}}\langle(\xi_{2},q_{2})\rangle^{s}{\lvert}\xi_{3}{\rvert}^{\frac{1}{8}}\langle(\xi_{3},q_{3})\rangle^{s}.$
An application of the dual version of (19), followed by Hölder’s inequality then yields
${\lVert}\partial_{x}(Q_{N_{1}}^{(\frac{1}{6})}(u_{1}u_{2})u_{3}){\rVert}_{X_{s,-\frac{1}{2}+2\varepsilon}}$
$\lesssim N_{1}^{\frac{3}{4}-2s}{\lVert}I_{x}^{\frac{1}{8}}(Q_{N_{1}}^{(\frac{1}{6})}(I_{x}^{\frac{1}{8}}J^{s}u_{1}I_{x}^{\frac{1}{8}}J^{s}u_{2})I_{x}^{\frac{1}{8}}J^{s}u_{3}){\rVert}_{X_{0,-\frac{1}{2}+2\varepsilon}}$
$\lesssim N_{1}^{\frac{3}{4}-2s}{\lVert}Q_{N_{1}}^{(\frac{1}{6})}(I_{x}^{\frac{1}{8}}J^{s}u_{1}I_{x}^{\frac{1}{8}}J^{s}u_{2})I_{x}^{\frac{1}{8}}J^{s}u_{3}{\rVert}_{L_{tx}^{\frac{4}{3}}L_{y}^{2}}$
$\leq N_{1}^{\frac{3}{4}-2s}{\lVert}Q_{N_{1}}^{(\frac{1}{6})}(I_{x}^{\frac{1}{8}}J^{s}u_{1}I_{x}^{\frac{1}{8}}J^{s}u_{2}){\rVert}_{L_{tx}^{2}L_{y}^{\infty}}{\lVert}I_{x}^{\frac{1}{8}}J^{s}u_{3}{\rVert}_{L_{tx}^{4}L_{y}^{2}},$
${\lVert}Q_{N_{1}}^{(\frac{1}{6})}(I_{x}^{\frac{1}{8}}J^{s}u_{1}I_{x}^{\frac{1}{8}}J^{s}u_{2}){\rVert}_{L_{y}^{\infty}}$ $\lesssim{\lVert}\chi_{{\{{\lvert}q{\rvert}\lesssim N_{1}^{\frac{1}{6}}\}}}\mathcal{F}_{y}(I_{x}^{\frac{1}{8}}J^{s}u_{1}I_{x}^{\frac{1}{8}}J^{s}u_{2}){\rVert}_{L_{q}^{1}}$
$\leq{\lVert}\chi_{{\{{\lvert}q{\rvert}\lesssim N_{1}^{\frac{1}{6}}\}}}{\rVert}_{L_{q}^{1}}{\lVert}\mathcal{F}_{y}(I_{x}^{\frac{1}{8}}J^{s}u_{1}I_{x}^{\frac{1}{8}}J^{s}u_{2}){\rVert}_{L_{q}^{\infty}}$
$\lesssim N_{1}^{\frac{1}{6}}{\lVert}I_{x}^{\frac{1}{8}}J^{s}u_{1}I_{x}^{\frac{1}{8}}J^{s}u_{2}{\rVert}_{L_{y}^{1}}.$
With this, we further obtain
${\lVert}\partial_{x}(Q_{N_{1}}^{(\frac{1}{6})}(u_{1}u_{2})u_{3}){\rVert}_{X_{s,-\frac{1}{2}+2\varepsilon}}$ $\lesssim N_{1}^{\frac{3}{4}+\frac{1}{6}-2s}{\lVert}I_{x}^{\frac{1}{8}}J^{s}u_{1}I_{x}^{\frac{1}{8}}J^{s}u_{2}{\rVert}_{L_{tx}^{2}L_{y}^{1}}{\lVert}I_{x}^{\frac{1}{8}}J^{s}u_{3}{\rVert}_{L_{tx}^{4}L_{y}^{2}}$
$\leq N_{1}^{\frac{3}{4}+\frac{1}{6}-2s}\prod_{i=1}^{3}{\lVert}I_{x}^{\frac{1}{8}}J^{s}u_{i}{\rVert}_{L_{tx}^{4}L_{y}^{2}},$
and three applications of (19) ultimately give
$...\lesssim N_{1}^{\frac{3}{4}+\frac{1}{6}-2s}\prod_{i=1}^{3}{\lVert}u_{i}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}}.$
The desired estimate (42) now follows upon performing a dyadic summation over all $N_{1}\sim N_{2}\sim N_{3}\gg 1$. This argument works provided that
$\frac{3}{4}+\frac{1}{6}-2s<0\Leftrightarrow\frac{11}{24}<s$
and $\varepsilon>0$ is chosen sufficiently small; this concludes the discussion of this subcase.
Under this assumption, the pointwise estimates ${\lvert}\xi_{0}{\rvert}\lesssim{\lvert}\xi_{1}+\xi_{2}{\rvert}$ and ${\lvert}\xi_{0}{\rvert}\lesssim{\lvert}\xi_{1}+\xi_{3}{\rvert}$ hold. Thus, if ${\lvert}3(\xi_{i}+\xi_{j})^{2}-(q_{i}+q_{j})^{2}{\rvert}\gtrsim{\lvert}\xi_{i}+\xi_{j}{\rvert}$ holds for at least one tuple $(i,j)\in{\{(1,2),(1,3)\}}$, then we may argue exactly as in case (ii.3.1), so that (30) and (33) lead to the desired multilinear estimate (42), valid for every $s>\frac{1}{4}$. Consequently, we may assume
${\lvert}3(\xi_{1}+\xi_{2})^{2}-(q_{1}+q_{2})^{2}{\rvert}\ll{\lvert}\xi_{1}+\xi_{2}{\rvert}\sim{\lvert}\xi_{3}{\rvert}$
and
${\lvert}3(\xi_{1}+\xi_{3})^{2}-(q_{1}+q_{3})^{2}{\rvert}\ll{\lvert}\xi_{1}+\xi_{3}{\rvert}\sim{\lvert}\xi_{3}{\rvert}$.
Let us now first consider the case where ${\lvert}\xi_{3}{\rvert}\ll{\lvert}(\xi_{1},q_{1}){\rvert}$. Then, invoking the general assumption (ii.3), we obtain
${\lvert}q_{1}{\rvert}\sim{\lvert}q_{2}{\rvert}\sim{\lvert}q_{3}{\rvert}\sim{\lvert}q_{0}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}$,
and in combination with the two constraints from the bilinear refinement, only the sign distributions $(q_{1},q_{2},q_{3},q_{0})=(\pm,\mp,\mp,\mp)$ remain possible. We once again distinguish two cases. In the case ${\lvert}\xi_{2}+\xi_{3}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, the resonance function $R$ in the form (43) provides a remedy: taking into account (ii.3), ${\lvert}\xi_{1}+\xi_{2}{\rvert}\sim{\lvert}\xi_{\text{max}}{\rvert}$, and ${\lvert}\xi_{1}+\xi_{3}{\rvert}\sim{\lvert}\xi_{\text{max}}{\rvert}$, we obtain
${\lvert}R{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}{\lvert}\xi_{\text{max}}{\rvert}^{2},$
and this - as already established in case (ii.3.1) - implies, at the very least, that an arbitrary $s>\frac{1}{3}$ can be reached in the multilinear estimate (42). Conversely, if ${\lvert}\xi_{2}+\xi_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, it follows that ${\lvert}\xi_{0}-\xi_{1}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, and, upon taking into account (ii.3), the magnitudes, and the signs of the $q_{i}$, one consequently obtains ${\lvert}q_{2}-q_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, ${\lvert}q_{0}+q_{1}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, and hence ${\lvert}q_{1}+q_{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$. Moreover, we have
${\lvert}\xi_{0}{\rvert}\leq{\lvert}\xi_{1}{\rvert}+{\lvert}\xi_{2}+\xi_{3}{\rvert}\leq{\lvert}\xi_{1}{\rvert}+\varepsilon_{0}{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}},$
and thus, ${\lvert}\xi_{0}{\rvert}\lesssim{\lvert}\xi_{1}{\rvert}$, so that the Airy $L_{tx}^{4}L_{y}^{2}$-estimate (19) in the previously carried out smallness argument can be applied a total of four times: Again, for dyadic numbers $N_{1},N_{2},N_{3}\in 2^{\mathbb{N}}$, we set
$u_{i}=P_{N_{i}}u_{i},\quad i\in\{1,2,3\},$
and obtain ${\lvert}q_{1}+q_{2}{\rvert}\ll N_{1}^{\frac{1}{3}}$ as well as
${\lvert}\xi_{0}{\rvert}\langle(\xi_{0},q_{0})\rangle^{s}\lesssim N_{1}^{\frac{1}{2}-2s}{\lvert}\xi_{0}{\rvert}^{\frac{1}{8}}\prod_{i=1}^{3}{\lvert}\xi_{i}{\rvert}^{\frac{1}{8}}\langle(\xi_{i},q_{i})\rangle^{s},$
which altogether - computing exactly as in case (ii.3.1) - yields
${\lVert}\partial_{x}(Q_{N_{1}}^{(\frac{1}{3})}(u_{1}u_{2})u_{3}){\rVert}_{X_{s,-\frac{1}{2}+2\varepsilon}}\lesssim N_{1}^{\frac{1}{2}+\frac{1}{3}-2s}\prod_{i=1}^{3}{\lVert}u_{i}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}}.$
A dyadic summation over all $N_{1}\sim N_{2}\sim N_{3}\gg 1$ now leads to the desired multilinear estimate (42), valid for all
$\frac{1}{2}+\frac{1}{3}-2s<0\Leftrightarrow\frac{5}{12}<s,$
provided $\varepsilon>0$ is chosen sufficiently small. With this, the discussion of the case ${\lvert}\xi_{3}{\rvert}\ll{\lvert}(\xi_{1},q_{1}){\rvert}$ is complete. We now consider the case ${\lvert}\xi_{3}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}$. If ${\lvert}\xi_{2}+\xi_{3}{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$ to begin with, no further steps are required, as we then have
${\lvert}R{\rvert}\gtrsim{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}{\lvert}\xi_{\text{max}}{\rvert}^{2}$
(observe (43), (ii.3), ${\lvert}\xi_{1}+\xi_{2}{\rvert}\sim{\lvert}\xi_{\text{max}}{\rvert}$, and ${\lvert}\xi_{1}+\xi_{3}{\rvert}\sim{\lvert}\xi_{\text{max}}{\rvert}$), and, as previously computed in case (ii.3.1), this ensures (at the very least) that all values $s>\frac{1}{3}$ can be realized within (42). We may thus assume ${\lvert}\xi_{2}+\xi_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, which directly implies ${\lvert}\xi_{0}{\rvert}\lesssim{\lvert}\xi_{1}{\rvert}$, and this fact will play a crucial role in a subsequent smallness argument. Now, if ${\lvert}q_{3}{\rvert}\gtrsim{\lvert}(\xi_{1},q_{1}){\rvert}$ were true, then, in conjunction with the active assumptions (ii.3), (ii.3.1), ${\lvert}\xi_{0}-\xi_{1}{\rvert}={\lvert}\xi_{2}+\xi_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, and the two constraints arising from the inapplicability of the bilinear refinement, we would obtain, on the one hand, the relations
${\lvert}q_{1}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert},\ {\lvert}{\lvert}q_{1}{\rvert}-{\lvert}q_{0}{\rvert}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}},\ \text{and}\ {\lvert}{\lvert}q_{2}{\rvert}-{\lvert}q_{3}{\rvert}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}},$
and on the other hand, the sign configurations $(q_{1},q_{2},q_{3},q_{0})=(\pm,\mp,\mp,\mp)$. These constraints, however, can only hold simultaneously in the case ${\lvert}q_{1}+q_{2}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, which would contradict ${\lvert}3(\xi_{1}+\xi_{2})^{2}-(q_{1}+q_{2})^{2}{\rvert}\ll{\lvert}\xi_{1}+\xi_{2}{\rvert}$. We must therefore have ${\lvert}q_{3}{\rvert}\ll{\lvert}(\xi_{1},q_{1}){\rvert}$ and consequently, by ${\lvert}\xi_{2}+\xi_{3}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$ and (ii.3), also ${\lvert}q_{2}{\rvert}\ll{\lvert}(\xi_{1},q_{1}){\rvert}$. Now, taking into account ${\lvert}q_{1}{\rvert}\sim{\lvert}(\xi_{1},q_{1}){\rvert}$, it follows that $q_{1}$ and $q_{0}$ must have the same sign, and together with (ii.3) and ${\lvert}\xi_{0}-\xi_{1}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$, this further implies that ${\lvert}q_{2}+q_{3}{\rvert}={\lvert}q_{0}-q_{1}{\rvert}\ll{\lvert}\xi_{0}{\rvert}^{\frac{1}{3}}$ must hold. After passing to dyadic pieces once again and taking into account ${\lvert}\xi_{0}{\rvert}\lesssim{\lvert}\xi_{1}{\rvert}$, we then obtain - as in the previous discussion - the estimate
${\lVert}\partial_{x}(u_{1}Q_{N_{1}}^{(\frac{1}{3})}(u_{2}u_{3})){\rVert}_{X_{s,-\frac{1}{2}+2\varepsilon}}\lesssim N_{1}^{\frac{1}{2}+\frac{1}{3}-2s}\prod_{i=1}^{3}{\lVert}u_{i}{\rVert}_{X_{s,\frac{1}{2}+\varepsilon}},$
which yields the desired trilinear estimate (42) after dyadic summation over all $N_{1}\sim N_{2}\sim N_{3}\gg 1$, provided that $\varepsilon>0$ is chosen sufficiently small and
$\frac{1}{2}+\frac{1}{3}-2s<0\Leftrightarrow\frac{5}{12}<s$
is satisfied. Thus, this subcase is also resolved.